Problem 6.29

One type of sunburn occurs on exposure to UV light of wavelength in the vicinity of 330 nm.

1)What is the energy of a photon of this wavelength?

2)What is the energy of a mole of these photons?

3)How many photons are in a 1.15 mJ burst of this radiation?

4)These UV photons can break chemical bonds in your skin to cause sunburn—a form of radiation damage. If the 330 nm radiation provides exactly the energy to break an average chemical bond in the skin, estimate the average energy of these bonds in kJ/mol.
1
Expert's answer
2016-10-20T18:52:09-0400
1) The energy of photon is E = hc/l = 6.62*10-34 J*sec * 3*108 m/sec / 3.3*10-7 m = 6.0*10-19 J (3.8 eV).
2) The energy of a mol of photons is Em = Na*E = 6.0*10-19 J * 6.02*10^23 mol-1 = 360 kJ/mol/
3) The number of photons is N = Eb/E = 1.15*10-3 J/ 6.0*10-19 J = 1.9*10^15.
4) The energy of an average bond is the same as the energy of a mole of photons - 360 kJ/mol.